# Dataset and Assessment of LLMs for Cultural Text Adaptation in Indian Context

### v1

@@ Abstract
Large language models (LLMs) are widely used in various tasks and applications. However, despite their wide capabilities, they are shown to lack cultural alignment (Ryan et al., 2024; AlKhamissi et al., 2024) and produce biased generations Naous et al. (2024) due to a lack of cultural knowledge and competence. Evaluation of LLMs for cultural awareness and alignment is particularly challenging due to the lack of proper evaluation metrics and unavailability of culturally grounded datasets representing the vast complexity of cultures at the regional and sub-regional levels. Existing datasets for culture specific items (CSIs) focus primarily on concepts at the regional level and may contain false positives. To address this issue, we introduce a novel CSI dataset for Indian culture, belonging to 17 cultural facets. The dataset comprises $\sim$8k cultural concepts from 36 sub-regions. To measure the cultural competence of LLMs on a cultural text adaptation task, we evaluate the adaptations using the CSIs created, LLM as Judge, and human evaluations from diverse socio-demographic region. Furthermore, we perform quantitative analysis demonstrating selective sub-regional coverage and surface-level adaptations across all considered LLMs. Our dataset is available here: https://huggingface.co/datasets/nlip/DIWALI, project webpage [^1] https://nlip-lab.github.io/nlip/publications/diwali/, and our codebase with model outputs can be found here: https://github.com/pramitsahoo/culture-evaluation.
@@ 1 Introduction
Figure 1: Comparison of the number of cultural concepts between CANDLE Nguyen et al. (2023) and Ours (DIWALI), highlighting differences in coverage of cultural concepts for five facets (clothing, drink, cuisine, rituals, and traditions) available in CANDLE (Domain = Country, Subject = India).
@@
Large language models (LLMs), despite having vast knowledge and being trained on extensive large data, are shown to lack cultural knowledge and competence Naous et al. (2024); Wang et al. (2024); Masoud et al. (2025). This is primarily attributed to highly biased pre-training data towards Western culture Li et al. (2024); Naous and Xu (2025); Li et al. (2025). Cultural text adaptation refers to modifying text from a general or specific culture to align with a particular target cultural group’s linguistic, social, and conceptual norms. Cultural adaptation has a wide range of applications, ranging from education Burstein et al. (2007), dialogue Singh et al. (2024), movie subtitling Ferreira (2024); El-Farahaty and Alwazna (2024), hospitality Radojević et al. (2024), advertisement Barthwal and Gupta (2011); Li (2024), cross-cultural communication Wilczewski et al. (2019), and storytelling Bhatt and Diaz (2024); Hamna et al. (2025). However, ensuring effective and emotionally resonating cultural adaptation is challenging, as it requires deep understanding and knowledge of social norms, subjective historical references, resonating events, and cultural nuances.
@@
With its vast linguistic and cultural diversity, India presents a unique challenge for culture-driven NLP systems. In particular, India houses 28 states and eight union territories [^2] https://knowindia.india.gov.in/states-uts/, each with multiple languages, dialects, traditions, and socio-cultural norms. Furthermore, cultural diversity exists at the state level and within sub-regions. For example, the state of Assam has communities like Bodo, Rabha, Karbi, and so on – each with distinct cultural norms, artifacts, and language [^3] https://assam.gov.in/about-us/391. Despite recent advances, existing LLMs struggle with cultural adaptation. Studies Naous et al. (2024) have shown that LLMs trained on English-centric corpora often fail to generate culturally appropriate content, especially for cultures from different parts of the world. In this work, we systematically evaluate multiple open-weight LLMs for cultural text adaptation on closed-text generation. In particular, we utilize existing text corpora from the education domain belonging to American culture. We rely on cultural concepts to assess the cultural competence of various LLMs for cultural text adaptation tasks. Culture Specific Items (CSIs) represent a particular culture’s concepts, objects, items, and customs. For example, Mekhela Chador (clothing), Pongal (food), and Darjeeling Tea (drink) are cultural concepts from India. Prior work, CANDLE Nguyen et al. (2023), has made foundational efforts to automatically extract culturally specific concepts from large web text for various cultures. In another work, DOSA Seth et al. (2024) builds community-driven cultural artifacts for a sub-regional culture belonging to India. However, these datasets have a limited representation of cultural concepts from India. Figure 1 shows the comparison of the created dataset with respect to existing concepts from the CANDLE [^4] https://candle.mpi-inf.mpg.de/ framework. To address this gap, we built a novel, large-scale dataset of CSIs for the Indian subregion. In particular, our dataset contains $\sim$8k cultural concepts from 17 facets. We evaluate using automatic CSI-based and human evaluation metrics to systematically assess the adaptation quality of various LLMs. Our dataset provides a valuable resource to the community for further research in the comprehensive and standardized assessment of cultural text adaptation. We define culture at the regional or country level, focusing on India.
@@ 2 Culture & Culture Specific Items
Culture is a multifaceted and complex construct. It means different things to different groups of people. It encompasses knowledge, beliefs, morals, values, and customs. Broadly, we can divide culture into material and non-material components. Material components include tangible elements such as food, dress, houses, and ornaments. In contrast, non-material components refer to symbols, ideas, beliefs, norms, values, morality, and attitudes that guide social behavior. For example, the practice of fasting during Navratris666a religious observance common in Northern India.. Culture is a complex concept and hard to define. As a result, the community has turned to using demographic and semantic proxies to approximately define culture Adilazuarda et al. (2024). Demographic proxies typically include geographical region, language, gender, race, religion, education, and ethnicity. In contrast, people associate semantic proxies with emotional expressions, food habits, social and political relations, actions, and naming conventions. In this work, we consider culture as a regional demographic proxy – with the geographical country India as a region and its states and union territories as sub-regions.
Newmark (2003) defined culture as “… the way of life and its manifestations that are peculiar to a community that uses particular languages as a means of expression.” This definition views culture as a collection of objects, customs, and traditions that hold meaning within a specific community. Newmark (2003) categorizes cultural manifestations into five collections or facets: (a) ecology, (b) material culture, (c) social culture, (d) organizations, customs, activities, procedures, concepts, and (e) gestures and habits. Culture Specific Items represent concepts, objects, items, and customs that hold significant cultural importance and relevance to a particular culture. It may vary across different subregions. Existing CSI datasets, such as CANDLE and DOSA, are limited in scope, covering only a subset of cultural items. CANDLE groups cultural common sense knowledge as a facet and concept. Each facet contains multiple concepts that are culturally relevant to a domain and subject. DOSA Seth et al. (2024) builds an India-specific cultural dataset by gamifying the process of data collection from 19 socio-diverse geographical locations, with a total of 615 social artifacts. Despite these foundational efforts, through an initial manual inspection, we observe that cultural concepts extracted by CANDLE are incorrectly mapped to a facet, for example: Light, Africa, Nepal under “Rituals” facet (Details in Appendix A).
@@ 4.2 Evaluation Dataset
We use GSM8k Cobbe et al. (2021) and its multilingual variant MGSM Shi et al. (2022) as our evaluation datasets. These benchmarks consist of grade school-level math word problems to evaluate the reasoning capabilities of LLMs. We select these datasets for two primary reasons. First, the educational domain is cognitively engaging and often reflects socio-cultural values, making it well-suited for evaluating cultural adaptation. Second, the problems usually contain culturally grounded entities such as names, food items, locations, and units, which can vary in relevance across cultures. For our experiments, we use the GSM8k (1,319 samples) and MGSM (250 samples) test sets, both originally situated in an American cultural context.
To assess robustness beyond math word problems, we additionally evaluate on two distinct domains: dialogue and story. Specifically, we sample 100 instances from DailyDialog Li et al. (2017) and 100 instances from ROCStories Mostafazadeh et al. (2016).
@@ 4.3 Models
We evaluate models from three families of large language models, with parameter sizes ranging from 1B to 9B: (a) Llama 2 7B Chat Touvron et al. (2023), (b) Llama 3.1 8B Instruct Grattafiori et al. (2024), (c) Llama 3.2 1B Instruct Grattafiori et al. (2024), (d) Llama 3.2 3B Instruct Grattafiori et al. (2024), (e) Mistral 7B Instruct Jiang et al. (2023), (f) Gemma 2 2B Instruct Team et al. (2024), and (g) Gemma 2 9B Instruct Team et al. (2024).
@@ 4.4.2 LLM as Judge
Motivated by recent progress in utilizing large language models to evaluate text generations Zheng et al. (2023). We conduct our evaluation using two state-of-the-art LLMs: Llama-3.1-8B-Instruct and Mistral 7B Instruct-v0.3. Both models are employed to assess the quality of culturally adapted text based on a detailed prompt and scoring criteria. In particular, we prompt LLMs to score adapted text in three dimensions: Cultural Relevance, Language Fluency, and Mathematical Integrity. We follow a Likert scale of 0 (very poor) to 5 (perfect).
@@ 5.3 Human Evaluation Results
To ensure consistency among annotators and the reliability of the scores. We measure the inter-annotator reliability using Cohen’s $\kappa$ Landis and Koch (1977) on the pair of human raters for each model. As shown in Table 5, there is acceptable agreement for models under the Llama and Gemma families. At the same time, evaluation scores for Mistral-7B-Instruct have low inter-annotator agreement.
@@ 6 Analyses
Sub-regional coverage: Despite their strong in-context capabilities and training on a large corpus, LLMs lack pluralistic Sorensen et al. (2024) alignment. This leads to the propagation of cultural ideas and views from their training data, causing a dominating effect while underrepresenting other cultures. To investigate this, we pose the following question: “Do LLMs, when prompted with the task of regional-level text adaptation, fairly represent all the sub-regional concepts?” To analyze this, we map each replaced concept with a sub-region (State or Union Territory) and plot a geographical heatmap for each facet. The heatmap for the food facet is shown in Figure 3, which reveals a significant bias in how concepts are adapted to represent a sub-region. In particular, food-related concepts adapted by Gemma-2-2B-Instruct primarily originate from Uttar Pradesh, Madhya Pradesh, Maharashtra, and Punjab, with a complete absence of adaptations from the Northeastern states. Similar sub-regional skews are observed for the rest of the facets. The heatmaps for the remaining 16 facets are presented in Appendix I.
@@
Surface vs. Deep level: Effective cultural adaptation must be faithful to the source text while emotionally resonating with the target audience by incorporating deep cultural nuances. Hershcovich et al. (2022) define culture along three axes: Aboutness, Objectives & Values, and Common ground. In particular, Aboutness refers to the contextual relevance of concepts across cultures. For instance, the discussion of “cricket match” might be less meaningful in cultures where cricket is less prevalent. Deeper adaptation connects culturally relevant events and scenarios. Therefore, cultural adaptation requires adapting from one culture to another by connecting relevant events and scenarios. For example, in the festival of “Durga puja” (event), people from Bengal might relate more to “pandal hopping” (scenario), thus impacting the aboutness of the text in context with culture. Such adaptations are at a deeper level and are more emotionally relatable.
@@ 7 Related Work
In this section, we describe the various works related to building specific cultural artifacts with a focus on Indian culture. CANDLE Nguyen et al. (2023) develops a large CSI collection by extracting cultural commonsense knowledge from web-scale corpora using an end-to-end method. It identifies candidate sentences using techniques such as string matching, named entity recognition, and a set of handcrafted lexico-syntactic rules. These candidates are then classified into cultural facets – such as geography, religion, or occupation – using zero-shot classification methods based on natural language inference. Finally, CANDLE clusters sentences containing similar facts using sentence embeddings and hierarchical agglomerative clustering, allowing it to produce coherent, representative summaries for each cluster. While this approach achieves broad coverage, its application to India is limited, consisting of only 650 concepts, and these entries focus on five facets: clothing, drinks, cuisines, rituals, and traditions.
In contrast, DOSA Seth et al. (2024) adopts a participatory research methodology by engaging 260 participants across 19 Indian states to compile a dataset of 615 artifacts. This method captures a broader range of cultural dimensions, including regional rituals, local customs, and specific art forms, providing richer insights into India’s sub-regional diversity. However, while providing depth, such participatory methods are difficult to scale. Our work addresses these gaps by constructing a comprehensive resource of cultural specific items, covering 17 facets across 36 sub-regions of India.
@@ 8 Usage of CSIs
Culture specific items (CSIs) are part of cultural knowledge bases and can potentially help elicit cultural knowledge in various applications, such as (a) cultural competence evaluation: existing methods on cultural competence evaluations of LLMs like Culture Bias Score Naous et al. (2024) and Edit level analysis Singh et al. (2024) rely on CSIs to measure the cultural biases of LLMs. Thus, having a large-scale, sub-regional CSIs for India would enable a more robust and nuanced evaluation of model competence across various downstream tasks, (b) cross-culture translation: while most machine translation systems are “culture agnostic,” recent work Conia et al. (2024); Yao et al. (2024); Zhang et al. (2024) has focused on culturally-aware translation. Having comprehensive CSI collection is essential for this task, as it allows models to identify and appropriately translate culture-specific entities for better cross-cultural translation, and (c) cultural adaptation: recent explorations in open text generation, such as writing a story Bhatt and Diaz (2024) and dialog-based Singh et al. (2024), have highlighted the need for cultural adaptation. CSIs can enable a better and holistic adaptation.

### v2

@@ Abstract
Large language models (LLMs) are widely used in various tasks and applications. However, despite their wide capabilities, they are shown to lack cultural alignment (ryan-etal-2024-unintended; alkhamissi-etal-2024-investigating) and produce biased generations naous-etal-2024-beer due to a lack of cultural knowledge and competence. Evaluation of LLMs for cultural awareness and alignment is particularly challenging due to the lack of proper evaluation metrics and unavailability of culturally grounded datasets representing the vast complexity of cultures at the regional and sub-regional levels. Existing datasets for culture specific items (CSIs) focus primarily on concepts at the regional level and may contain false positives. To address this issue, we introduce a novel CSI dataset for Indian culture, belonging to 17 cultural facets. The dataset comprises $\sim$8k cultural concepts from 36 sub-regions. To measure the cultural competence of LLMs on a cultural text adaptation task, we evaluate the adaptations using the CSIs created, LLM as Judge, and human evaluations from diverse socio-demographic region. Furthermore, we perform quantitative analysis demonstrating selective sub-regional coverage and surface-level adaptations across all considered LLMs. Our dataset is available here: https://huggingface.co/datasets/nlip/DIWALI, project webpage [^1] https://nlip-lab.github.io/nlip/publications/diwali/, and our codebase with model outputs can be found here: https://github.com/pramitsahoo/culture-evaluation.
@@ 1 Introduction
Figure 1: Comparison of the number of cultural concepts between CANDLE 10.1145/3543507.3583535 and Ours (DIWALI), highlighting differences in coverage of cultural concepts for five facets (clothing, drink, cuisine, rituals, and traditions) available in CANDLE (Domain = Country, Subject = India).
@@
Large language models (LLMs), despite having vast knowledge and being trained on extensive large data, are shown to lack cultural knowledge and competence naous-etal-2024-beer; wang-etal-2024-countries; masoud-etal-2025-cultural. This is primarily attributed to highly biased pre-training data towards Western culture NEURIPS2024_9a16935b; naous-xu-2025-origin; liattributing. Cultural text adaptation refers to modifying text from a general or specific culture to align with a particular target cultural group’s linguistic, social, and conceptual norms. Cultural adaptation has a wide range of applications, ranging from education burstein-etal-2007-automated, dialogue singh-etal-2024-translating, movie subtitling Ana_Luisa_Ferreira_2024; doi:10.1177/21582440241258293, hospitality Radojević_Stanišić_Stanić_2024, advertisement Barthwal_Gupta_2011; jopem13, cross-cultural communication WILCZEWSKI2019362, and storytelling bhatt-diaz-2024-extrinsic; 10.1145/3715335.3735478. However, ensuring effective and emotionally resonating cultural adaptation is challenging, as it requires deep understanding and knowledge of social norms, subjective historical references, resonating events, and cultural nuances.
@@
With its vast linguistic and cultural diversity, India presents a unique challenge for culture-driven NLP systems. In particular, India houses 28 states and eight union territories [^2] https://knowindia.india.gov.in/states-uts/, each with multiple languages, dialects, traditions, and socio-cultural norms. Furthermore, cultural diversity exists at the state level and within sub-regions. For example, the state of Assam has communities like Bodo, Rabha, Karbi, and so on – each with distinct cultural norms, artifacts, and language [^3] https://assam.gov.in/about-us/391. Despite recent advances, existing LLMs struggle with cultural adaptation. Studies naous-etal-2024-beer have shown that LLMs trained on English-centric corpora often fail to generate culturally appropriate content, especially for cultures from different parts of the world. In this work, we systematically evaluate multiple open-weight LLMs for cultural text adaptation on closed-text generation. In particular, we utilize existing text corpora from the education domain belonging to American culture. We rely on cultural concepts to assess the cultural competence of various LLMs for cultural text adaptation tasks. Culture Specific Items (CSIs) represent a particular culture’s concepts, objects, items, and customs. For example, Mekhela Chador (clothing), Pongal (food), and Darjeeling Tea (drink) are cultural concepts from India. Prior work, CANDLE 10.1145/3543507.3583535, has made foundational efforts to automatically extract culturally specific concepts from large web text for various cultures. In another work, DOSA seth-etal-2024-dosa builds community-driven cultural artifacts for a sub-regional culture belonging to India. However, these datasets have a limited representation of cultural concepts from India. Figure 1 shows the comparison of the created dataset with respect to existing concepts from the CANDLE [^4] https://candle.mpi-inf.mpg.de/ framework. To address this gap, we built a novel, large-scale dataset of CSIs for the Indian subregion. In particular, our dataset contains $\sim$8k cultural concepts from 17 facets. We evaluate using automatic CSI-based and human evaluation metrics to systematically assess the adaptation quality of various LLMs. Our dataset provides a valuable resource to the community for further research in the comprehensive and standardized assessment of cultural text adaptation. We define culture at the regional or country level, focusing on India.
@@ 2 Culture & Culture Specific Items
Culture is a multifaceted and complex construct. It means different things to different groups of people. It encompasses knowledge, beliefs, morals, values, and customs. Broadly, we can divide culture into material and non-material components. Material components include tangible elements such as food, dress, houses, and ornaments. In contrast, non-material components refer to symbols, ideas, beliefs, norms, values, morality, and attitudes that guide social behavior. For example, the practice of fasting during Navratris666a religious observance common in Northern India.. Culture is a complex concept and hard to define. As a result, the community has turned to using demographic and semantic proxies to approximately define culture adilazuarda-etal-2024-towards. Demographic proxies typically include geographical region, language, gender, race, religion, education, and ethnicity. In contrast, people associate semantic proxies with emotional expressions, food habits, social and political relations, actions, and naming conventions. In this work, we consider culture as a regional demographic proxy – with the geographical country India as a region and its states and union territories as sub-regions.
newmark2003textbook defined culture as “… the way of life and its manifestations that are peculiar to a community that uses particular languages as a means of expression.” This definition views culture as a collection of objects, customs, and traditions that hold meaning within a specific community. newmark2003textbook categorizes cultural manifestations into five collections or facets: (a) ecology, (b) material culture, (c) social culture, (d) organizations, customs, activities, procedures, concepts, and (e) gestures and habits. Culture Specific Items represent concepts, objects, items, and customs that hold significant cultural importance and relevance to a particular culture. It may vary across different subregions. Existing CSI datasets, such as CANDLE and DOSA, are limited in scope, covering only a subset of cultural items. CANDLE groups cultural common sense knowledge as a facet and concept. Each facet contains multiple concepts that are culturally relevant to a domain and subject. DOSA seth-etal-2024-dosa builds an India-specific cultural dataset by gamifying the process of data collection from 19 socio-diverse geographical locations, with a total of 615 social artifacts. Despite these foundational efforts, through an initial manual inspection, we observe that cultural concepts extracted by CANDLE are incorrectly mapped to a facet, for example: Light, Africa, Nepal under “Rituals” facet (Details in Appendix A).
@@ 4.2 Evaluation Dataset
We use GSM8k cobbe2021trainingverifierssolvemath and its multilingual variant MGSM shi2022languagemodelsmultilingualchainofthought as our evaluation datasets. These benchmarks consist of grade school-level math word problems to evaluate the reasoning capabilities of LLMs. We select these datasets for two primary reasons. First, the educational domain is cognitively engaging and often reflects socio-cultural values, making it well-suited for evaluating cultural adaptation. Second, the problems usually contain culturally grounded entities such as names, food items, locations, and units, which can vary in relevance across cultures. For our experiments, we use the GSM8k (1,319 samples) and MGSM (250 samples) test sets, both originally situated in an American cultural context.
To assess robustness beyond math word problems, we additionally evaluate on two distinct domains: dialogue and story. Specifically, we sample 100 instances from DailyDialog li-etal-2017-dailydialog and 100 instances from ROCStories mostafazadeh-etal-2016-corpus.
@@ 4.3 Models
We evaluate models from three families of large language models, with parameter sizes ranging from 1B to 9B: (a) Llama 2 7B Chat touvron2023Llama2openfoundation, (b) Llama 3.1 8B Instruct grattafiori2024Llama3herdmodels, (c) Llama 3.2 1B Instruct grattafiori2024Llama3herdmodels, (d) Llama 3.2 3B Instruct grattafiori2024Llama3herdmodels, (e) Mistral 7B Instruct jiang2023mistral7b, (f) Gemma 2 2B Instruct gemmateam2024gemma2improvingopen, and (g) Gemma 2 9B Instruct gemmateam2024gemma2improvingopen.
@@ 4.4.2 LLM as Judge
Motivated by recent progress in utilizing large language models to evaluate text generations zheng2023judging. We conduct our evaluation using two state-of-the-art LLMs: Llama-3.1-8B-Instruct and Mistral 7B Instruct-v0.3. Both models are employed to assess the quality of culturally adapted text based on a detailed prompt and scoring criteria. In particular, we prompt LLMs to score adapted text in three dimensions: Cultural Relevance, Language Fluency, and Mathematical Integrity. We follow a Likert scale of 0 (very poor) to 5 (perfect).
@@ 5.3 Human Evaluation Results
To ensure consistency among annotators and the reliability of the scores. We measure the inter-annotator reliability using Cohen’s $\kappa$ 8d20e0b8-89d8-3d65-bcf5-8c19d56ec4ab on the pair of human raters for each model. As shown in Table 5, there is acceptable agreement for models under the Llama and Gemma families. At the same time, evaluation scores for Mistral-7B-Instruct have low inter-annotator agreement.
@@ 6 Analyses
Sub-regional coverage: Despite their strong in-context capabilities and training on a large corpus, LLMs lack pluralistic 10.5555/3692070.3693952 alignment. This leads to the propagation of cultural ideas and views from their training data, causing a dominating effect while underrepresenting other cultures. To investigate this, we pose the following question: “Do LLMs, when prompted with the task of regional-level text adaptation, fairly represent all the sub-regional concepts?” To analyze this, we map each replaced concept with a sub-region (State or Union Territory) and plot a geographical heatmap for each facet. The heatmap for the food facet is shown in Figure 3, which reveals a significant bias in how concepts are adapted to represent a sub-region. In particular, food-related concepts adapted by Gemma-2-2B-Instruct primarily originate from Uttar Pradesh, Madhya Pradesh, Maharashtra, and Punjab, with a complete absence of adaptations from the Northeastern states. Similar sub-regional skews are observed for the rest of the facets. The heatmaps for the remaining 16 facets are presented in Appendix I.
@@
Surface vs. Deep level: Effective cultural adaptation must be faithful to the source text while emotionally resonating with the target audience by incorporating deep cultural nuances. hershcovich-etal-2022-challenges define culture along three axes: Aboutness, Objectives & Values, and Common ground. In particular, Aboutness refers to the contextual relevance of concepts across cultures. For instance, the discussion of “cricket match” might be less meaningful in cultures where cricket is less prevalent. Deeper adaptation connects culturally relevant events and scenarios. Therefore, cultural adaptation requires adapting from one culture to another by connecting relevant events and scenarios. For example, in the festival of “Durga puja” (event), people from Bengal might relate more to “pandal hopping” (scenario), thus impacting the aboutness of the text in context with culture. Such adaptations are at a deeper level and are more emotionally relatable.
@@ 7 Related Work
In this section, we describe the various works related to building specific cultural artifacts with a focus on Indian culture. CANDLE 10.1145/3543507.3583535 develops a large CSI collection by extracting cultural commonsense knowledge from web-scale corpora using an end-to-end method. It identifies candidate sentences using techniques such as string matching, named entity recognition, and a set of handcrafted lexico-syntactic rules. These candidates are then classified into cultural facets – such as geography, religion, or occupation – using zero-shot classification methods based on natural language inference. Finally, CANDLE clusters sentences containing similar facts using sentence embeddings and hierarchical agglomerative clustering, allowing it to produce coherent, representative summaries for each cluster. While this approach achieves broad coverage, its application to India is limited, consisting of only 650 concepts, and these entries focus on five facets: clothing, drinks, cuisines, rituals, and traditions.
In contrast, DOSA seth-etal-2024-dosa adopts a participatory research methodology by engaging 260 participants across 19 Indian states to compile a dataset of 615 artifacts. This method captures a broader range of cultural dimensions, including regional rituals, local customs, and specific art forms, providing richer insights into India’s sub-regional diversity. However, while providing depth, such participatory methods are difficult to scale. Our work addresses these gaps by constructing a comprehensive resource of cultural specific items, covering 17 facets across 36 sub-regions of India.
@@ 8 Usage of CSIs
Culture specific items (CSIs) are part of cultural knowledge bases and can potentially help elicit cultural knowledge in various applications, such as (a) cultural competence evaluation: existing methods on cultural competence evaluations of LLMs like Culture Bias Score naous-etal-2024-beer and Edit level analysis singh-etal-2024-translating rely on CSIs to measure the cultural biases of LLMs. Thus, having a large-scale, sub-regional CSIs for India would enable a more robust and nuanced evaluation of model competence across various downstream tasks, (b) cross-culture translation: while most machine translation systems are “culture agnostic,” recent work cultural-translation; yao-etal-2024-benchmarking; zhang-etal-2024-cultural has focused on culturally-aware translation. Having comprehensive CSI collection is essential for this task, as it allows models to identify and appropriately translate culture-specific entities for better cross-cultural translation, and (c) cultural adaptation: recent explorations in open text generation, such as writing a story bhatt-diaz-2024-extrinsic and dialog-based singh-etal-2024-translating, have highlighted the need for cultural adaptation. CSIs can enable a better and holistic adaptation.